OpenStudy (anonymous):

WILL MEDAL AND GIVE DANK MEMES FOR ANSWER The equation of line CD is y = −2x − 2. Write an equation of a line parallel to line CD in slope-intercept form that contains point (4, 5). y = −2x + 13 y = negative 1 over 2 x + 7 y = negative 1 over 2 x + 3 y = −2x − 3

OpenStudy (anonymous):

You here for the dank memes?

OpenStudy (anonymous):

after you help me

OpenStudy (welshfella):

not sure what dank nemes are!

OpenStudy (anonymous):

I will show you after you help

OpenStudy (anonymous):

I think the answer is a

OpenStudy (welshfella):

ok well a line that is parallel to y = -2x -2 will have the same slope The -2x indicates that the slope is -2.

OpenStudy (anonymous):

Yes

OpenStudy (welshfella):

To find if the point (4,5) is on the line you plug in x = 4 and y = 5 into the equation and see if it fits.

OpenStudy (anonymous):

so the equation would be 2(x-4) = y-5

OpenStudy (anonymous):

distribute it and get 2x-8=y-5

OpenStudy (welshfella):

the equATION WOULD BE y = -2x + a where a is a number to be found

OpenStudy (anonymous):

opposite operations and get -2x +8=y-5

OpenStudy (anonymous):

i forgot that the slope was -2

OpenStudy (anonymous):

not 2

OpenStudy (welshfella):

substituting 5 = -2(4) + 13 LHS = 5 RHS = -8+13 = 5 so A is the correct equatuon

OpenStudy (anonymous):

yeah

OpenStudy (welshfella):

oh ok